When 1.0000g of a compound containing C, H, and N is combusted, 2.6667 g of CO2 and 5455g of H2o is produced...
a) determine the g of C in this compound
b) determine the g of H in this compound
c) determine the g of N in this compound
d)determine the empirical and the molecular formula of this compound
1
Expert's answer
2012-10-08T10:02:51-0400
The compound has formula C(x)H(y)N(z). The combustion reaction is
C(x)H(y)N(z) + O2 = xCO2 + (y/2)H2O + (z/2)N2
x to y ratio is
(2.6667/44):(2*0.5455/18)=
0.06061/0.06061=1 : 1
Thus, the compound formula is
CHN(z). The Molecular mass of the compound is
(12 + 1 + 14z)
For x=1, the empirical formula is
CHN4.
The molecular formula is C2H2N8 -
5,5'-bistetrazole.